OpenStudy (goformit100):

The diameter of a circle equals the radius of another circle. The ratio of their areas is (A) 1 : 2 (B) 1 : 3 (C) 1 : 4 (D) 1 : 8 (E) 1 : 16

OpenStudy (goformit100):

@ganeshie8

OpenStudy (anonymous):

\[\frac{A_1}{A_2} = \frac{\pi r_1^2}{\pi r_2^2}\] \[2r_1=r_2\]

OpenStudy (goformit100):

Thanks

OpenStudy (anonymous):

\[\frac{A_1}{A_2} = \frac{r_1^2}{(2r)_1^2}\]